# Limit-case admissibility for positive infinite-dimensional systems

### v1

@@ 5 Applications to perturbation results
Let $U$ be a Banach space and $Z$ be an ordered Banach space. In Section 4, the property that an operator $T\in\mathcal{L}(U,Z)$ maps the unit ball of $U$ into an order-bounded subset of $Z$ played an important role in obtaining sufficient conditions for admissibility. In this appendix, we give a characterisation and, for positive $T$, a number of sufficient conditions for this property. We note in passing that if $Z$ is even a Banach lattice, the slightly weaker property that $T$ maps the unit ball of $U$ into an order-bounded subset of $Z^{\prime\prime}$ is called majorizing [47, Proposition IV.3.4].
@@
Let $U$ be a Banach space and let $Z$ be an ordered Banach space with a normal cone. The following conditions are equivalent for an operator $T\in\mathcal{L}(U,Z)$.
@@
If $Z$ is, in addition, a KB-space, then (i) and (ii) are also equivalent to each of the following conditions:
@@
Proof A.2.
First, assume that (ii) holds and let $u\in B_{U}$. Since $\widetilde{X}$ has a unit, there exists – according to [22, Proposition 2.11] – an element $e\in\widetilde{X}$ and $\epsilon>0$ such that the implication
$\left\lVert x\right\rVert\leq\epsilon\Rightarrow x\leq e$
holds for all $x\in\widetilde{X}$. In particular, taking $\lambda=\epsilon^{-1}\left\lVert T_{1}\right\rVert$, it follows that $T_{1}u\in[-\lambda e,\lambda e]$; where we have used $\left\lVert u\right\rVert\leq 1$. Employing the positivity of $T_{2}$, it follows that $T=T_{2}T_{1}$ maps the unit ball of $U$ into an order-bounded subset of $Z$.
“(i) $\Rightarrow$ (ii)”: The observation about ordered bounded subsets made prior to the proposition guarantees the existence of $e\in Z_{+}$ such that $T(B_{U})\subseteq[-e,e]$. On the other hand, because the cone of $Z$ is normal, $\widetilde{X}:=Z_{e}$ is an ordered Banach space with a unit. The above argument allows us to infer that $\operatorname{Rg}T\subseteq\widetilde{X}$. Whence, $T$ factors through $\widetilde{X}$ via the canonical embedding $\widetilde{X}\hookrightarrow Z$ (which is positive).
Finally, suppose that $Z$ is a KB-space, then $Z_{+}$ is a face of $(Z^{\prime\prime})_{+}$. The equivalence of (i)-(iv) now follows by [47, Proposition IV.3.4].
@@
Let $U$ and $Z$ be ordered Banach spaces. A positive operator $T\in\mathcal{L}(U,Z)$ maps the unit ball of $U$ into an order-bounded subset of $Z$ in each of the following cases:
@@
    The closed unit ball of $U$ is upwards directed and every increasing norm-bounded net in $Z_{+}$ is norm-convergent.
Proof A.4.
@@
(b) Let $S=B_{U}\cap U_{+}$. Since $B_{U}$ is upwards directed, for each $u_{1},u_{2}\in S$, there exists $u\in S$ such that $u_{1},u_{2}\leq u$. On the other hand, the assumption on $Z$ implies that the net $(Tu)_{u\in S}$ is norm convergent in $Z$; here we’ve used the positivity of $T$. It follows that $T$ maps the unit ball of $U$ into an order-bounded subset of $Z$.
Appendix B Lower estimate for operators and a characterisation of AL-spaces
Let $E$ be a Banach lattice and let $A:\operatorname{dom}A\subseteq E\to E$ be a resolvent positive operator, i.e., all sufficiently large real numbers $\lambda$ are in the resolvent set of $A$ and satisfy $R(\lambda,A)\geq 0$. In particular, this is satisfied if $A$ generates a positive $C_{0}$-semigroup on $E$. Consider the following condition:
(B.1) There exist numbers $\begin{split}\text{There exist numbers }\lambda>&s(A)\text{ and }c>0\text{ such that }A\\ \text{ satisfies }\left\lVert R(\lambda,A)f\right\rVert&\geq c\left\lVert f\right\rVert\text{ for every }f\in E_{+}.\end{split}$
This condition occurs in a recent paper on infinite-dimensional positive systems [19, Theorems 2.1] where it is shown to have rather strong consequences for admissibility of control operators. Earlier, the condition (B.1) was used to show generation results [5, Theorem 2.5].
The condition (B.1) is, of course, satisfied in the trivial case where $A$ is bounded. More interestingly, if $E$ is an $L^{1}$-space and the semigroup generated by $A$ is stochastic – i.e., positive and norm preserving on $E_{+}$ – then it is also easy to see that (B.1) is satisfied. However, it is not clear at all how to find an unbounded operator on an $L^{p}$-space for $p>1$ that satisfies (B.1). In this appendix we justify this theoretically: in the important special case where $A$ has compact resolvent, the condition (B.1) implies that $E$ is – up to an equivalent renorming – an $L^{1}$-space (Corollary B.3).
Let us recall the following terminology again: An AL-space is a Banach lattice whose norm is additive on the positive cone. Every $L^{1}$-space is an AL-space and conversely, every AL-space can be shown to be isomorphic (as a Banach lattice) to an $L^{1}$-space (over a possible non-$\sigma$-finite measure space). Compare also Table 1 in the introduction. We first show a characterisation of AL-spaces, up to equivalent norms. For related geometric results, see [36, Proposition 3.11] and [2, Section 2].
@@ Appendix B Lower estimate for operators and a characterisation of AL-spaces
For each Banach lattice $E$, the following are equivalent:
@@
    There exists an equivalent norm on $E$ which turns $E$ into an AL-space.
@@
    No net in the positive unit sphere of $E$ converges weakly to 00.
Proof B.2.
“(i) $\Rightarrow$ (ii)”: This can easily seen by testing against the norm functional on an AL-space.
“(ii) $\Rightarrow$ (i)”: Let (ii) be true. Let $\mathcal{F}^{\prime}$ denote the set of all non-empty finite subsets of the positive unit sphere in $E^{\prime}$. Assume for a moment that the following property ()(*)( ∗ ) holds: For each number $\varepsilon>0$ and each $F^{\prime}\in\mathcal{F}^{\prime}$, there exists a vector $x_{\varepsilon,F^{\prime}}$ in the positive unit sphere of $E$ such that $\langle x^{\prime},x_{\varepsilon,F^{\prime}}\rangle<\varepsilon$ for all $x^{\prime}\in F^{\prime}$. The relation $\preceq$ on $(0,\infty)\times\mathcal{F}^{\prime}$ that is given by $(\varepsilon_{1},F_{1}^{\prime})\preceq(\varepsilon_{2},F_{2}^{\prime})$ if and only if $\varepsilon_{1}\geq\varepsilon_{2}$ and $F_{1}^{\prime}\subseteq F_{2}^{\prime}$, turns $(0,\infty)\times\mathcal{F}^{\prime}$ into a directed set, and the net $(x_{\varepsilon,F^{\prime}})_{(\varepsilon,F^{\prime})\in(0,\infty)\times\mathcal{F}^{\prime}}$ is contained in the positive unit sphere of $E$ and converges weakly to 00. This contradicts (ii), so we conclude that the property ()(*)( ∗ ) does not hold.
Hence, there exists a number $\varepsilon>0$ and a set $F^{\prime}\in\mathcal{F}^{\prime}$ such that for all $x$ in the positive unit sphere of $E$ one has $\langle x^{\prime},x\rangle\geq\varepsilon$ for at least one element $x^{\prime}\in F^{\prime}$. So if we set $z^{\prime}:=\sum_{x^{\prime}\in F^{\prime}}x^{\prime}$, then we obtain $\langle z^{\prime},x\rangle\geq\varepsilon\left\lVert x\right\rVert$ for each $x\in E_{+}$. Hence, $x\mapsto\langle z^{\prime},\left\lvert x\right\rvert\rangle$ is an equivalent lattice norm on $E$, which is clearly additive on the positive cone.
@@
Let $E$ be a Banach lattice and let $T:E\to E$ be a compact linear operator. If there exists a number $c>0$ such that $\left\lVert Tx\right\rVert\geq c\left\lVert x\right\rVert$ for all $x\in E_{+}$, then there exists an equivalent norm on $E$ which turns $E$ into an AL-space.
Proof B.4.
Assume that there is no equivalent norm on $E$ which turns $E$ into an AL-space. Then there exists, by Theorem B.1, a net $(x_{j})$ in the positive unit sphere of $E$ which converges weakly to 00. Since $T$ is compact, it turns weakly convergent nets into norm convergent ones, so $(Tx_{j})$ converges to 00 in norm. However, one has $\left\lVert Tx_{j}\right\rVert\geq c$ for each index $j$ by assumption; this is a contradiction.
@@
Let $E$ be a Banach lattice and let $A:\operatorname{dom}A\subseteq E\to E$ be a resolvent positive linear operator. If $A$ has compact resolvent and satisfies (B.1), then there exists an equivalent norm on $E$ which turns $E$ into an AL-space.
Acknowledgements
The first author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 523942381. The article was initiated during a pleasant visit of the first author to the third author at the Tampere University, Finland in autumn 2023. The first author is indebted to COST Action 18232 for financial support for this visit. The third author was funded by the Research Council of Finland grant number 349002.

### v4

@@ Appendix A Order-bounded images of the unit ball
In Section 4, the property that an operator $T\in\mathcal{L}(U,Z)$ maps the unit ball of a Banach space $U$ into an order-bounded subset of an ordered Banach space $Z$ played an important role in obtaining sufficient conditions for admissibility. In this appendix, we give a characterisation and, for positive $T$, a number of sufficient conditions for this property. Note that if $Z$ is even a Banach lattice, the weaker property that $T$ maps the unit ball of $U$ into an order-bounded subset of $Z^{\prime\prime}$ is called majorizing [53, Proposition IV.3.4].
@@
Let $U$ be a Banach space, let $Z$ be an ordered Banach space with a normal cone, and let $T\in\mathcal{L}(U,Z)$. The following are equivalent.
@@
If $Z$ is even a KB-space, then (i) and (ii) are also equivalent to each of the following:
@@
Proof.
“(ii) $\Rightarrow$ (i)”: Assume that (ii) holds. Since $\widetilde{X}$ has a unit, there exists – according to [27, Proposition 2.11] – an element $e\in\widetilde{X}$ and $\varepsilon>0$ such that the implication
$\left\lVert x\right\rVert\leq\varepsilon\quad\Rightarrow\quad x\leq e$
holds for all $x\in\widetilde{X}$. In particular, taking $\lambda=\varepsilon^{-1}\left\lVert T_{1}\right\rVert$, it follows that $T_{1}u\in[-\lambda e,\lambda e]$ for every $u\in B_{U}$. Employing the positivity of $T_{2}$, it follows that $T=T_{2}T_{1}$ maps the unit ball of $U$ into an order-bounded subset of $Z$.
“(i) $\Rightarrow$ (ii)”: The observation about order bounded subsets made prior to the proposition guarantees the existence of $e\in Z_{+}$ such that $T(B_{U})\subseteq[-e,e]$. On the other hand, because the cone of $Z$ is normal, $\widetilde{X}:=Z_{e}$ is an ordered Banach space with a unit. The above argument allows us to infer that $\operatorname{Rg}T\subseteq\widetilde{X}$. Whence, $T$ factors through $\widetilde{X}$ via the canonical embedding $\widetilde{X}\hookrightarrow Z$ (which is positive).
Finally, suppose that $Z$ is a KB-space. Then $Z_{+}$ is a face of $(Z^{\prime\prime})_{+}$ according to [63, Theorem 7.1]. The equivalence of (i)–(iv) now follows from [53, Proposition IV.3.4]. ∎
@@
Let $U$ and $Z$ be ordered Banach spaces. Then $T\in\mathcal{L}(U,Z)_{+}$ maps the unit ball of $U$ into an order-bounded subset of $Z$ in the following cases:
@@
    The open unit ball of $U$ is upwards directed and every increasing norm-bounded net in $Z_{+}$ is norm-convergent.
Proof.
@@
(b) Let $C$ denote the open unit ball of $U$. Since $C$ is directed, $(Tu)_{u\in C}$ is an increasing and norm bounded net in $Z$ and thus, by assumption, norm converges to an element $z\in Z$. Hence, $TC\subseteq[-z,z]$ and since order intervals are closed, we conclude that $T$ also maps the closed unit ball $B_{U}=\overline{C}$ into $[-z,z]$. ∎
Appendix B Lower resolvent bounds and a characterisation of AL-spaces
Let $X$ be a Banach lattice and let $A:\operatorname{dom}A\subseteq X\to X$ be a resolvent positive operator, i.e., all sufficiently large real numbers $\lambda$ are in the resolvent set of $A$ and satisfy $R(\lambda,A)\geq 0$. In particular, this is satisfied if $A$ generates a positive $C_{0}$-semigroup on $X$. Consider the following condition:
There exist numbers $\begin{split}\text{There exist numbers }\lambda>&s(A)\text{ and }c>0\text{ such that }A\\ \text{ satisfies }\left\lVert R(\lambda,A)f\right\rVert&\geq c\left\lVert f\right\rVert\text{ for every }f\in X_{+}.\end{split}$ (B.1)
This condition occurs in recent papers on infinite-dimensional positive systems, see [22, Theorems 2.1] and [23, Theorem 1]) and was shown to have rather strong consequences for admissibility of control operators. Earlier (B.1) was used to show generation results [5, Theorem 2.5].
The condition (B.1) is, of course, satisfied in the trivial case where $A$ is bounded. More interestingly, if $X$ is an $L^{1}$-space and the semigroup generated by $A$ is stochastic – i.e., positive and norm preserving on $X_{+}$ – then it is also easy to see that (B.1) is satisfied. However, it is not clear at all how to find an unbounded operator on an $L^{p}$-space for $p>1$ that satisfies (B.1). In this appendix we justify this theoretically: in the important special case where $A$ has compact resolvent, the condition (B.1) implies that $X$ is – up to an equivalent renorming – an $L^{1}$-space (Corollary B.2).
Let us recall the following terminology again: An AL-space is a Banach lattice whose norm is additive on the positive cone. Every $L^{1}$-space is an AL-space and conversely, every AL-space can be shown to be isomorphic (as a Banach lattice) to an $L^{1}$-space (over a possible non-$\sigma$-finite measure space), cf. Table 1. We first show a characterisation of AL-spaces, up to equivalent norms. For related geometric results, see [42, Proposition 3.11] and [2, Section 2].
@@ Appendix B Lower resolvent bounds and a characterisation of AL-spaces
For a Banach lattice $X$, the following are equivalent:
@@
    There exists an equivalent norm on $X$ which turns $X$ into an AL-space.
@@
    No net in the positive unit sphere of $X$ converges weakly to 00.
Proof.
“(i) $\Rightarrow$ (ii)”: This can easily be seen by testing against the norm functional on the AL-space.
“(ii) $\Rightarrow$ (i)”: Let $\mathcal{F}^{\prime}$ denote the set of all non-empty finite subsets of the positive unit sphere in $X^{\prime}$. Assume for a moment that the following property ()(*)( ∗ ) holds: For each number $\varepsilon>0$ and each $F^{\prime}\in\mathcal{F}^{\prime}$, there exists a vector $x_{\varepsilon,F^{\prime}}$ in the positive unit sphere of $X$ such that $\langle x^{\prime},x_{\varepsilon,F^{\prime}}\rangle<\varepsilon$ for all $x^{\prime}\in F^{\prime}$. The relation $\preceq$ on $(0,\infty)\times\mathcal{F}^{\prime}$ given by $(\varepsilon_{1},F_{1}^{\prime})\preceq(\varepsilon_{2},F_{2}^{\prime})$ if and only if $\varepsilon_{1}\geq\varepsilon_{2}$ and $F_{1}^{\prime}\subseteq F_{2}^{\prime}$, turns $(0,\infty)\times\mathcal{F}^{\prime}$ into a directed set. The net $(x_{\varepsilon,F^{\prime}})_{(\varepsilon,F^{\prime})\in(0,\infty)\times\mathcal{F}^{\prime}}$ in the positive unit sphere of $X$ converges weakly to 00, contradicting (ii), so the property ()(*)( ∗ ) cannot hold.
Hence, there exists a number $\varepsilon>0$ and a set $F^{\prime}\in\mathcal{F}^{\prime}$ such that for all $x$ in the positive unit sphere of $E$ one has $\langle x^{\prime},x\rangle\geq\varepsilon$ for some $x^{\prime}\in F^{\prime}$. Setting $z^{\prime}:=\sum_{x^{\prime}\in F^{\prime}}x^{\prime}$, we obtain $\langle z^{\prime},x\rangle\geq\varepsilon\left\lVert x\right\rVert$ for each $x\in X_{+}$. Hence, $x\mapsto\langle z^{\prime},\left\lvert x\right\rvert\rangle$ is an equivalent lattice norm on $X$, which is clearly additive on the positive cone. ∎
@@
Let $X$ be a Banach lattice and let $T:X\to X$ be a compact linear operator. If there exists a number $c>0$ such that $\left\lVert Tx\right\rVert\geq c\left\lVert x\right\rVert$ for all $x\in X_{+}$, then there exists an equivalent norm on $X$ which turns $X$ into an AL-space.
Proof.
If the conclusion is false, then by Theorem B.1, there is a net $(x_{j})$ in the positive unit sphere of $X$ converging weakly to 00. From compactness of $T$, we deduce that $(Tx_{j})$ converges to 00 in norm, contradicting that $\left\lVert Tx_{j}\right\rVert\geq c$ for each index $j$. ∎
@@
Let $X$ be a Banach lattice and let $A:\operatorname{dom}A\subseteq X\to X$ be a resolvent positive linear operator. If $A$ has compact resolvent and satisfies (B.1), then there exists an equivalent norm on $X$ which turns $X$ into an AL-space.